### v1

# World In Your Hands: A Large-Scale and Open-source Ecosystem for Learning Human-centric Manipulation in the Wild

Abstract
Large-scale pre-training is fundamental for generalization in language and vision models, but data for dexterous hand manipulation remains limited in scale and diversity, hindering policy generalization. Limited scenario diversity, misaligned modalities, and insufficient benchmarking constrain current human manipulation datasets. To address these gaps, we introduce World In Your Hands (WiYH), a large-scale open-source ecosystem for human-centric manipulation learning. WiYH includes (1) the Oracle Suite, a wearable data collection kit with an auto-labeling pipeline for accurate motion capture; (2) the WiYH Dataset, featuring over 1,000 hours of multi-modal manipulation data across hundreds of skills in diverse real-world scenarios; and (3) extensive annotations and benchmarks supporting tasks from perception to action. Furthermore, experiments based on the WiYH ecosystem show that integrating WiYH’s human-centric data significantly enhances the generalization and robustness of dexterous hand policies in tabletop manipulation tasks. We believe that World In Your Hands will bring new insights into human-centric data collection and policy learning to the community.
1 Introduction
| Dataset | Hours | Clips | Wild | Skills | Cam. Calib | RGB | Tactile | Action | Depth | Mask | Instruct. | Reason |
| Ego4D [grauman2022ego4d] | 3,670 | - | ✓ | ✓ | ✗ | ✓ | ✗ | 2D | ✗ | ✗ | ✓ | ✓ |
| HOI4D [liu2022hoi4d] | 44.4 | 4k | ✓ | ✓ | ✓ | ✓ | ✗ | 2D | ✓ | ✓ | ✗ | ✗ |
| HoloAssist [wang2023holoassist] | 166 | 2.22k | ✗ | ✗ | ✗ | ✓ | ✗ | - | ✓ | ✗ | ✓ | ✓ |
| Ego-Exo4D [grauman2024ego] | 1,286 | 5.04k | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✗ | ✓ | ✓ | ✓ |
| CaptainCook4D [peddi2024captaincook4d] | 94.5 | 384 | ✓ | ✓ | ✗ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | ✓ |
| HOT3D [banerjee2025hot3d] | 13.9 | 425 | ✗ | ✗ | ✓ | ✓ | ✗ | 3D | ✗ | ✗ | ✗ | ✗ |
| HO-Cap [wang2024ho] | - | 64 | ✗ | ✗ | ✓ | ✓ | ✗ | 3D | ✓ | ✓ | ✗ | ✗ |
| Ego-ViD-5M [wang2024egovid] | - | 5M | ✓ | ✗ | ✓ | ✓ | ✗ | ✗ | ✗ | ✗ | ✓ | ✗ |
| EgoDex [hoque2025egodex] | 829 | 338k | ✗ | ✓ | ✗ | ✓ | ✗ | 3D | ✗ | ✓ | ✓ | ✗ |
| WiYH (Ours) | 1045 | 125.4k | ✓ | ✓ | ✓ | ✓ | ✓ | 3D | ✓ | ✓ | ✓ | ✓ |
Large-scale pre-training has been established as a cornerstone for achieving generalization in Large Language Models [brown2020language, chowdhery2023palm, achiam2023gpt, touvron2023llama, yang2025qwen3], Vision-Language Models [radford2021learning, li2022blip, alayrac2022flamingo, li2023blip, li2024llava] and Vision-Language-Action Models [brohan2022rt, zitkovich2023rt, kim2024openvla, black2410pi0, wen2024diffusion, zhou2025vision]. However, datasets for dexterous hand manipulation remain far smaller and less diverse than language datasets, limiting progress in learning robust manipulation policies.
To address the problem of data scarcity, human-centric data collection and learning have attracted growing interest [hsieh2025dexman, zhu2025learning, routray2025vipra, zhu2025let, kareer2025egomimic]. Some approaches [chao2021dexycb, kwon2021h2o, banerjee2025hot3d] capture human hand data during object manipulation using devices such as VR systems and smart glasses, while others [grauman2022ego4d, xu2021videoclip, wang2021actionclip, he2024egovm] leverage foundation models to extract human actions from unlabeled videos of human manipulation available on the internet.
Despite these advances, current human manipulation datasets still face three major limitations, as shown in Table 1 that hinder their usefulness for dexterous manipulation research: (1) Limited Scenario Diversity: Many datasets are collected in constrained laboratory environments, lacking the diversity and complexity of real-world settings. (2) Inadequate Alignment: Existing collections often lack one or more key modalities, such as the absence of fine-grained language instructions, precise 3D hand poses, or object poses, resulting in misalignment across vision, language, and action. (3) Insufficient Benchmarking: Most datasets do not support comprehensive evaluation benchmarks that span the entire pipeline of dexterous manipulation, including scene perception, task planning, and action.
To bridge this gap and investigate in-depth how human-centric data can advance dexterous manipulation, we introduce World In Your Hands (WiYH), a large-scale open-source ecosystem for learning human-centric manipulation in the wild, as shown in Figure 1. Specifically, the WiYH ecosystem comprises three core components:
The Oracle Suite and Auto-labeling Pipeline. We introduce a lightweight and flexible wearable data collection kit, the Oracle Suite, designed to capture data without impeding the operator’s natural movements. An integrated auto-labeling pipeline enables markerless self-localization and centimeter-level motion accuracy without requiring external third-person tracking.
The WiYH Dataset: A Large-Scale Multi-modal Human-centric Dataset. We present the WiYH dataset, a large-scale multi-modal resource comprising over 1,000 hours of manipulation sequences across hundreds of skills, captured in 10 diverse industrial and daily-life scenarios such as factories, supermarkets, hotels, and restaurants. The dataset includes rich sensory streams: multi-view images, camera calibration parameters, depth maps, 3D hand poses, and 6D wrist trajectories.
Diverse Annotations and Comprehensive Benchmarks. Building upon the WiYH dataset, we provide extensive annotations, including approximately 400 hours of atomic-level action instructions and 100k instances of vision-language data for task reasoning, affordance understanding, and object localization. These annotations support a comprehensive set of benchmarks—spanning visual language understanding, world model evaluation, and vision-language-action tasks—designed to assess capabilities from scene perception to embodied operation holistically.
Based on the WiYH ecosystem, we study how human-centric data empowers dexterous manipulation learning in tabletop manipulation tasks. We train models on mixtures of Oracle Suite demonstrations and other manipulation data and evaluate the task success rate across varied environments and object layouts. Our experiments demonstrate that human-centric data substantially improves the generalization and robustness of dexterous hand policies.
[Limitation]. All manipulation tasks are evaluated in real-world environments. To increase accessibility and reproducibility, we are developing high-fidelity simulation environments and plan to release them in a future version of the dataset and benchmarks.
2 World In Your Hands Ecosystem
2.1 Oracle Suite: Human-centric Data Collection Suite
To enable large-scale, high-fidelity, human-centric multimodal data acquisition, we developed Oracle Suite, a low-cost wearable data collection system designed for human-centric embodied AI research. Specifically, as shown in Figure 2, Oracle Suite comprises a hardware system for acquiring multi-view RGB, IMU-based localization, and tactile information, alongside an autolabeling module that achieves high-precision motion capture through the fusion of IR, RGB, and IMU localization.
Hardware System Overview. Oracle Suite adopts a modular architecture integrating three primary hardware components: H-HPVHive, H-Gloves, and H-Backpack. This modular design enables flexible deployment across diverse environments, without being restricted to controlled indoor labs.
The Oracle Suite introduces major optimizations in ergonomics, sensing, and reliability:
H-FPVHive: It supports chest-mounted configuration, and integrates two fisheye cameras, two pinhole cameras, and four infrared lenses. All cameras are hardware synchronized. The infrared lenses are used for better locating the relative poses of H-Gloves.
H-Gloves: Each glove contains 6 IMUs, 5 pressure sensors, and an on-board MCU. The gloves can capture the hand motions and tactile. Moreover, it also contains 3 fisheye cameras for visual tracking and observation.
H-Backpack: It provides data storage, computation, and power supply. The computation is hosted by NVIDIA Orin.
Autolabeling Pipeline. As shown in the right part of Figure 2, the Autolabeling Pipeline comprises both online and offline algorithms. It fuses multi-modal sensor data (IR + IMU + RGB) to ensure positioning accuracy in in-the-wild scenarios. Specifically, the online localization module performs lightweight coarse positioning using visual-inertial odometry (VIO) algorithms and IR detection, while the offline stage refines these coarse estimates through visual Structure-from-Motion (SfM) algorithms. The pipeline ultimately outputs centimeter-level 6D poses for both wrists.
Data Production Efficiency. The Oracle Suite pipeline follows a collection–annotation–inspection–archiving workflow for scalable data production. A single collector operating 5 hours per day generates approximately 1.8 TB of valid multimodal data. At scale, 100 collectors over 30 days yield roughly 5.4 PB of raw data (stored in cloud repositories). Efficiency is further improved through standardized capture protocols and automated cloud-based annotation.
2.2 Data Annotation Pipeline
Data collection: Practitioners wore the Oracle Suite in each scene to record natural manipulation behavior. Each recording session lasted approximately 10 minutes and included start and end event markers. The collected signals comprise tactile signals from pressure sensors, binocular fisheye images from the left and right wrist mounts, first-person binocular fisheye images, left-hand and right-hand IMU localization data, and the corresponding camera calibration parameters. Samples of data annotations are shown in Figure 4.
Data annotation: The annotation pipeline comprises three parts: subtask (atomic task) annotation, perception annotation, and vision-language annotation.
Subtask annotation: To construct a large-scale, high-quality set of atomic tasks, we designed a four-stage hybrid annotation pipeline consisting of data processing, task definition, annotation consolidation, and post-processing:
(1) Data processing: First-person fisheye videos are segmented into 30-second clips, which serve as the basic units for semantic annotation by the video-language module;
(2) Task definition (manual annotation): The video-language module parses hierarchical operational semantics from the first-person videos and hierarchically decomposes an operation sequence into high-level tasks and atomic subtask units. Each task unit contains a preset action described in natural language and a manually annotated target object.
(3) Data checking (manual review): Outputs from the video-language module are subject to iterative manual review; any incorrect annotations are reprocessed before proceeding to the consolidation stage.
(4) Post-processing: Annotated preset-action and target-object names are assembled into complete operation instructions and are further augmented using a large language model (LLM). After subtask annotation, the raw data are organized into a hierarchical format of tasks composed of multiple atomic subtasks.
Perception annotation: To provide richer perceptual information, we pre-annotated scene elements using visual foundation models and then performed manual selection and correction. The resulting perception annotations include:
(1) Segmentation masks: instance masks are generated by applying SAM to human-provided prompt points;
(2) Depth maps: scene depth is estimated from first-person binocular images using the FoundationStereo model.
Vision-language annotation: The vision-language annotations were produced through manual curation and comprise the following tasks:
(1) Spatial Referring: For each subtask, we extract a key frame near the action start that contains the operator’s object of attention and in which the target object is not occluded by the hand, body, or other objects. Based on the subtask’s target object, spatial relational terms are added to uniquely refer to a single target region in the image.
(2) Subtask Prediction: We select tasks that contain 4–5 subtasks. The next subtask following the current one is used as the correct option. Nine incorrect options are generated by combining scene object information and the task’s preset actions; annotators then select the three incorrect options that are the most logically plausible. Additionally, Qwen-2.5VL-72B was used to generate auxiliary “thinking” data for next-action prediction, which were subsequently corrected by human annotators;
(3) Completion Verification: We select tasks containing 4-5 subtasks and randomly trim the tail of the video by 20%, 25%, 30%, or 35% of frames. The task description and the truncated video are provided as input, and human annotators label whether the video depicts a completed task.
Through the above procedures, we obtain rich annotations covering scene understanding and task planning, which constitute the foundation of a unified benchmark.
2.3 Data Statistics
Figure 3 summarizes key statistics of our WiYH dataset. We first analyze the distribution of tasks across scenes. As shown in Figure 3(a), WiYH contains a diverse set of daily environments, including banquet, laundry, logistics, hotel, department, office, supermarket, industry, cleaning, and candlelight settings. The Sankey diagram illustrates the total accumulated duration of each task within its corresponding scene.
Scenes such as supermarkets and laundry include long-horizon activities with substantial execution time, while hotels and candlelight contain shorter but highly structured routines. This diversity in task–scene composition highlights the wide operational range covered by WiYH. Action-level temporal characteristics are shown in Figure 3(b). For each scene, we compute the average execution duration of fine-grained actions. The dataset exhibits substantial variation across action categories, from short, instantaneous operations such as wipe and insert, to long-horizon manipulation such as smooth out or foldlay. These temporal differences reflect the heterogeneous complexity of real-world household and industrial workflows.
Figure 3(c) reports the distribution of annotation types. WiYH includes rich multimodal and task-level supervision, consisting of RGB, depth, action, instruction, reasoning, mask, tactile, and calibration annotations. Among them, RGB and calibration frames appear most frequently, while reasoning and tactile annotations, though smaller in quantity, provide high-level semantic cues crucial for complex manipulation and planning tasks.
To better understand the manipulation space in WiYH, Figure 3(d–e) visualizes word clouds of the annotated target objects and skills. The dataset covers a broad range of manipulable objects—including clothes, linens, cartons, cups, and tableware—spanning deformable, rigid, and articulated categories. Likewise, the skill vocabulary encompasses diverse manipulation primitives such as take, place, rotate, unfold, align, push, and search, representing both low-level motor actions and higher-level task strategies. These distributions further demonstrate that WiYH captures a wide spectrum of real-world manipulation behaviors.
3 Benchmark and Applications
3.1 Human-centric Vision-Language (HVL)
In this section, we construct three evaluation tasks based on the vision-language annotations introduced in Section 2. These tasks are designed to evaluate whether existing Multimodal Large Language Models (MLLMs) can achieve a genuine understanding of human-centric manipulation tasks in highly generalized in-the-wild environments.
We define the following three evaluation tasks and test several closed-source and open-source MLLMs, including GPT-4o [hurst2024gpt], Doubao-seed [bai2024seed], and Qwen [yang2025qwen3]. Detailed experimental setups and data statistics are provided in the supplementary material. (1) Spatial Referring (SR): Given a human manipulation image and a description of the spatial relationships among multiple objects in the scene, the model is required to output the pixel coordinates of a specified region. Performance is measured by the probability that the predicted point lies within the target area. (2) Subtask Prediction (SP): Given a piece of human manipulation video, the current task description, and four candidate options for the next action, the model must select the correct subsequent action. (3) Completion Verification (CV): Given a piece of human manipulation video and the current task description, the model needs to determine whether the task has been completed.
As shown in Table 2, Qwen3-VL-Plus, which was pre-trained on spatial understanding data, achieves the best performance on the Spatial Referring task. However, scores of all MLLMs are below 0.5, indicating that current MLLMs still lack sufficient capability in interpreting fine-grained spatial relationships in manipulation scenarios. On the Completion Verification task, all models perform slightly above 0.5, suggesting that their ability to accurately judge task completion in such contexts remains limited. These experimental results collectively demonstrate that general-purpose, vanilla VLLMs still fall short in embodied tasks. Our WiYH dataset, with its rich annotations of human activities and spatial relationships, presents a promising foundation for future embodied pre-training.
| Model | SR (%) | SP (%) | CV (%) |
| GPT-4o [openai2024gpt4o] | |||
| Qwen3-VL-Plus [qwen2025vl] | |||
| Doubao-Seed-1.6-Vision [doubao2024seed] | |||
| Qwen3-VL-4B-Instruct [qwen2025vl_4b] |
3.2 Human-centric World Model (HWM)
| CogVideo [hong2022cogvideo] | DynamiCrafter [xing2024dynamicrafter] | |||
| Metric | w/o WiYH | w/ WiYH | w/o WiYH | w/ WiYH |
| Consistency | 88.2 (+7.6) | 91.5 (+6.2) | ||
| Smoothness | 98.4 (+4.9) | 98.6 (+1.2) | ||
| Dynamic | 78.5 (+15.6) | 89.6 (+14.8) | ||
| Quality | 67.6 (+7.5) | 70.6 (+7.1) | ||
3.2.1 Language-conditioned Video Generation
This section evaluates the contribution of our proposed WiYH dataset to enhancing video generation performance in human-centric manipulation tasks. We fine-tuned two baseline models, CogVideo [hong2022cogvideo] and DynamiCrafter [xing2024dynamicrafter], initialized with their original weights, using WiYH under the image-and-text-to-video setting. During fine-tuning, all input videos were resized to 480p resolution.
As shown in Figure 5, models fine-tuned with WiYH generate egocentric videos with improved spatial and temporal coherence, as well as stronger alignment with both the input image and action instructions. For example, in the “Wipe the cabinet” task, the cabinet’s orientation and the towel’s color are reproduced more accurately. In contrast, videos generated by non-fine-tuned models exhibit noticeable object fragmentation and distortion, along with discontinuous and physically implausible hand–object interactions. Notably, after fine-tuning, the complex motions of non-rigid objects in the generated videos appear more realistic and natural. These improvements demonstrate that WiYH substantially enhances a model’s ability to capture real-world dynamics.
Quantitative results are summarized in Table 3. We use four metrics from VBench [huang2024vbench]—consistency, smoothness, dynamic (motion intensity), and quality—to assess motion plausibility, temporal coherence, and visual fidelity. After fine-tuning with the WiYH dataset, both baseline models show clear improvements across all metrics, with the most significant gains observed in motion intensity and temporal consistency.
3.2.2 Gaussian Splatting 4D Reconstruction
This subsection evaluates the dataset’s utility for supervised end-to-end 4D reconstruction using Gaussian Splatting (GS). By providing high-quality RGB images, depth maps, and camera poses, WiYH supports shape and motion modeling for dynamic scenes. We extracted geometry using MegaSAM [megasam], tracked pixel trajectories of dynamic objects with TAPIR [tapir], and performed scene reconstruction with Shape of Motion [shapeofmotion]. The successful application of our dataset in spatial perception tasks like 4D Gaussian Splatting (4DGS) demonstrates the critical role of its rich annotations, particularly depth information, in enhancing spatial understanding and representation.
As shown in Table 4, photometric metrics remain relatively consistent, whereas geometric metrics vary notably across tasks. These differences highlight the inherent challenges in reconstructing dynamic scenes, where accurate geometry prediction is particularly complex in motion-rich environments. This observation suggests promising directions for future research in 4D reconstruction for robotic manipulation, particularly methods that improve geometric understanding in complex dynamic scenes.
Figure 6 presents the 4D reconstruction results at various time steps, showcasing the dataset’s potential for dynamic scene modeling. By capturing high-quality dynamic scenes, WiYH can help leverage Gaussian Splatting (GS) as an effective 3D tokenizer, advance spatial understanding, and enrich Real2Sim pipelines.
| Scene | Task | Photometry | Geometry | |||
| PSNR | SSIM | LPIPS | A.R. | |||
| 1 | Pouring Wine | 20.15 | 0.815 | 0.206 | 1.04 | 52.01 |
| 2 | Packing Clothes | 20.02 | 0.656 | 0.423 | 2.89 | 17.83 |
| 3 | Packing Cargo | 18.83 | 0.749 | 0.552 | 3.52 | 13.79 |
3.3 Human-centric VLA Training (HVT)
In this section, we evaluate several Vision-Language-Action (VLA) algorithms trained on the WiYH dataset to explore whether its visual-action-language tri-modal support enables effective training of vision-language-action-aligned manipulation policies.
Specifically, we use GR00T [bjorck2025gr00t] and OpenVLA-OFT [kim2025fine] as baseline algorithms, training each to predict the future 1.6-second motion trajectory (16 points at 10 Hz) of the human wrist from the dataset. We sample 10% of the data from each scenario for training and assemble a test set of 400 non-overlapping episodes. Each baseline model is trained for 30,000 steps with a total batch size of 128 on 8 NVIDIA A100 GPUs.
As shown in Table 5, training with human-centric data consistently reduces the error in tracking the ground-truth trajectories on the test set across both mainstream VLA frameworks. These results demonstrate WiYH’s effectiveness for training VLA algorithms. We believe the WiYH dataset can serve as a foundational resource for validating scaling laws in human-centric VLA pre-training.
| Model | Mean Error (cm) | ||
| 0 Steps | 15k Steps | 30k Steps | |
| OpenVLA-OFT [kim2025fine] | 4.75 | 1.28 | 1.12 |
| GR00T [bjorck2025gr00t] | 4.54 | 1.31 | 1.07 |
4 Cross-Embodiment Analysis
In this section, we conduct co-training experiments using cross-embodiment data to investigate the effectiveness of the human video data.
4.1 Implementation Details
We set up two scenarios: a single-object scene and a multi-object cluttered scene. In this experiment, we employ a co-training policy using data collected by an improved version of DexUMI [xu2025dexumi] and from Human Video demonstrations. Figure 8 shows representative scene examples and the data collection setup. For UMI, data collection always started from a limited number of initial poses. The improved UMI system is capable of capturing trajectories with millimeter-level accuracy and was exclusively used to collect high-precision manipulation data in the single-object scene, resulting in fine-grained motion trajectories. In contrast, Human Video contains only manipulation sequences from multi-object cluttered scenes, providing data that spans a larger action space, exhibits a broader distribution of motions, and covers a greater diversity of object observations.
4.2 Results and In-depth Discussion
In the single-object scenario, the robot arm begins execution from a random initial pose during testing, which differs from the initial pose used in the UMI dataset. In the multi-object cluttered scenario, we primarily evaluate the performance of the co-training strategy under conditions of varying scene height, changes in the target object’s pose, and increased occlusion. The configurations of these test scenarios and qualitative experimental results are illustrated in Figure 7.
Quantitative results are summarized in Table 6. (1) In the single-object scenario, co-training the policy with both UMI data and Human Video data leads to an improvement in the success rate of more than 13%. Although the motion precision in the human data is limited, its diverse action distribution enhances the policy’s generalization capability to unseen initial states. (2) Policies trained solely on UMI data exhibit almost no generalization ability in complex scenarios; however, after co-training with human data, the SR increases by 52%. This demonstrates that incorporating human data significantly improves the policy’s ability to interpret complex scenes. Therefore, the co-training approach effectively introduces both the action domain and observation domain of human data into the robot dataset, thereby enhancing the generalization capability of the learned policy.
| Scenario | Data Sources and Amounts | SR (%) | |
| UMI (clips) | Human Video (clips) | ||
| Single-object Scene | 200 | 0 | 20 |
| 200 | 800 | 23.3 (+3.3) | |
| 500 | 0 | 43.3 | |
| 500 | 800 | 56.7 (+13.4) | |
| Cluttered Scene | 200 | 0 | 0.0 |
| 200 | 800 | 32.0 (+32.0) | |
| 500 | 0 | 8.0 | |
| 500 | 800 | 60.0 (+52.0) | |
5 Conclusion
This paper presents the WiYH ecosystem, a cornerstone for embodied intelligence research in real-world environments. We make three key contributions: the Oracle Suite, a wearable system enabling in-the-wild data capture with integrated markerless auto-labeling; the large-scale WiYH Dataset, offering diverse, skilled human manipulation sequences with rich multi-modal streams; and a suite of Comprehensive Benchmarks, supported by extensive atomic-action and vision-language annotations, for holistic evaluation from perception to physical interaction. Our benchmarks reveal that current models still lack the fine-grained spatial and causal reasoning required for robust embodied operation. The WiYH dataset, with its unique scale, and rich annotations, is designed to bridge this gap. We believe this work provides an essential foundation for developing more capable and generalizable embodied agents.
Appendix A World in Your Hand (WiYH) Dataset
A.1 Detailed Task-Skill Counting Statistics
We further supplement the visualization of the number of skills corresponding to each task. Due to space limitations, we report statistics based on only 25% of the dataset, as shown in Figure. 16. The full visualization analysis will be updated subsequently on the dataset website.
A.2 Additional annotation examples visualizations
We provide additional annotation visualizations for the remaining 6 scenes, as shown in Figure 15.
A.2.1 Data Annotation Details.
Raw videos are structured into a three-tier hierarchy: Scene - Task - Subtask by Annotators. Each subtask corresponds to the minimal object-centered action unit or a temporally cohesive action sequence. (1) Subtask Annotation:
Actions: Drawn from predefined operation sets (e.g., PICK, PLACE_INTO, SEARCH).
Target Object: Semantic structure describing action objects and positional relationships (e.g., [object, container]).
Temporal Bounds: Precise start/end timestamps delimiting the execution segment in the video.
Status: Labeled as success, failure, irrelevant to denote execution outcome.
(2) Subtask Prediction chain-of-thought (CoT): With the subtask annotation, we select tasks of moderate length (containing 4-10 subtasks). We annotate them with Chain-of-Thought data, providing reasoning that decomposes the task into several logically connected subtasks. The annotation pipeline for CoT is illustrated in Figure 10. Initially, we provide the model (Qwen2.5-VL) with both the current subtask information (video and text) and the ground-truth text of the next subtask. The model is then prompted to perform posterior attribution, explaining why the specified next subtask logically follows. Subsequently, the attribution is appended to the input and presented again to Qwen2.5-VL, which synthesizes it into a structured causal reasoning process. Finally, the CoT segment is extracted and used to prompt Qwen-3, which is instructed to predict an answer. This predicted answer is matched against the ground-truth answer to validate the completeness and soundness of the CoT labeling.
Appendix B Benchmark Details
B.1 Human-centric Vision-Language
B.1.1 Benchmark Task Descriptions.
The Human-centric Vision-Language Benchmark in Section 3.1 is constructed by sampling raw data from WiYH, followed by manual annotation, including Spatial Referring, Subtask Prediction, and Completion Verification, with distributions illustrated in Figure 9. (1) The Spatial Referring is designed as a spatial region grounding form: given relevant image frames, annotators delineate ROI areas using polygons according to the given questions. It evaluates spatial understanding by prompting the model to output a set of points representing the spatial region referenced in the question. The primary metric is the proportion of these points that fall within the ground-truth polygon. (2) The Subtask Prediction adopts a multiple-choice question format. Using Qwen2.5-VL-72B, we first generate 9 negative candidate options based on video segments and ground truth, spanning action, object, and orientation dimensions. Annotators then manually select the top three hard negative samples, which, together with the correct answer, form the final candidate set. (3) The Completion Verification is designed as binary questions: given a video segment randomly truncated from the end, annotators determine whether the subtask has been completed. And this evaluation metric is the selection accuracy rate.
B.1.2 Benchmark Tasks Visualization.
Examples of the three tasks are shown in Figure 11.
B.2 Human-centric World Model
B.2.1 Language-conditioned Video Generation
Baseline. We validate the effectiveness of our WiYH dataset using video generation baselines with different architectures, including transformer (CogVideo [hong2022cogvideo]) and U-Net (DynamiCrafter [xing2024dynamicrafter]).
Evaluation Metrics. Following VBench [huang2024vbench], the evaluation metrics are as follows:
Consistency. To evaluate the temporal consistency of the target manipulation subjects, we calculate the DINO [caron2021emerging] feature similarity across frames. For assessing the temporal consistency of the background scenes, we compute the CLIP [radford2021learning] feature similarity across frames. The overall consistency score is defined as the arithmetic mean of the subject consistency and the background consistency.
Smoothness. To assess the temporal continuity of the generated videos, we employ a motion-consistency metric based on the AMT [li2023amt] video frame interpolation model. Given a generated video with frames , we first remove all odd-indexed frames to obtain . The AMT model is then used to interpolate the missing frames, producing . Finally, we compute the mean absolute error between the interpolated and original frames, where a larger value indicates smoother and more temporally coherent motion.
Dynamic. To prevent the model from producing static videos in pursuit of higher temporal consistency scores, we employ RAFT [teed2020raft] to estimate optical flow strengths between consecutive frames of each generated video. The average of the largest 5% optical flows is used to determine whether the video is static. The final dynamic score is calculated by measuring the proportion of non-static videos generated by the model.
Quality. To assess the visual quality of generated video frames, we employ the MUSIQ [ke2021musiq] image quality predictor trained on the SPAQ [fang2020perceptual] dataset. This metric quantifies common distortions such as overexposure, noise, and blur, with a higher score indicating fewer artifacts.
Implementation Details. During fine-tuning, all input videos are resized to 480p resolution. For CogVideo, we use the pre-trained image-to-video version with a temporal length of 49, a learning rate of 2e-5, and a batch size of 16. For Dynamicrafter, we leverage the model pre-trained at 512 resolution, setting the temporal length to 48, the learning rate to 1e-5, and the batch size to 16.
B.2.2 Gaussian Splatting 4D Reconstruction
Evaluation Metrics. In terms of metrics of 4D reconstruction, we evaluate the RGB rendering quality using PSNR, SSIM, and LPIPS. For PSNR and SSIM, higher is better, while for LPIPS, lower is better. For geometry accuracy from rendered depth maps, we use A.R. (lower is better) and (higher is better).
PSNR (Peak Signal-to-Noise Ratio),
where is the reference image, is the rendered image, and is the maximum possible pixel value. A higher PSNR indicates that the rendered image is closer to the ground truth.
SSIM (Structural Similarity Index Measure).
where , , and denote the mean, variance, and covariance of image patches and , respectively. SSIM evaluates perceptual similarity in terms of luminance, contrast, and structure; values closer to indicate better similarity.
LPIPS (Learned Perceptual Image Patch Similarity). This metric computes the perceptual distance between rendered and ground-truth images by extracting deep features from a pretrained CNN and measuring the distance in that feature space. Lower LPIPS values indicate that two images are more perceptually similar.
AbsRel (Absolute Relative Error).
where is the predicted depth at the pixel and is the corresponding ground-truth depth. A lower value denotes a more accurate geometry.
(Threshold Accuracy / Delta1).
which measures the percentage of predicted depth values that are within a factor of of the ground-truth depth. A higher value indicates better accuracy.
Implementation Details. During optimization, all input videos are processed at a resolution of 960×720. We represent each scene with 40k dynamic Gaussians and 100k static Gaussians and use B = 10 shared SE(3) motion bases for all points. We optimize the model with Adam, running 1k iterations for the initialization stage and 500 training epochs, with learning rates of for Gaussian means and SE(3) bases, for scales, for rotations, and for colors, opacities, and motion coefficients. Training on a 300-frame sequence takes about 2 hours on a single A100 GPU and supports real-time rendering at around 140 FPS.
B.3 Details of Cross-Embodiment Analysis.
In this section, we present the details of our cross-embodiment experiments in Section 4, Cross-Embodiment Analysis, of the main text. As shown in Figure 14, we demonstrate 11 task variants under two scenarios: Single-Object Scene and Multi-Object Cluttered Scene. Figures 12 and 13 illustrate the success rates and failure causes for each task in these two scenarios, respectively.
As shown in Figures 12 and 13, on most tasks, the incorporation of human data leads to a general increase in success rate. In the Single-Object Scene, the diverse action distribution of human data enhanced the policy’s generalization capability to unseen initial states, enabling successful grasping at different positions or over long distances. In the Multi-Object Cluttered Scene, policies trained solely on UMI data demonstrated almost no generalization ability in complex scenarios. When factors such as object height or layout changed, policies trained only on UMI data failed, whereas the inclusion of human data gradually enabled success by effectively introducing both the action and observation domains of human data into the robot dataset.
Furthermore, the consistently low success rates across all methods in certain tasks highlight important directions for future research: how to further improve the collection precision of human data and how to design efficient algorithms that can better transfer knowledge from human data to dexterous manipulation tasks.
Appendix C Ethics and Privacy Statement
Our work involves large-scale data collection of real-world human manipulation activities. We took a number of measures to ensure responsible data handling, privacy protection, and ethical compliance.
Data Collection and Consent. All human subjects participating in data collection were recruited voluntarily and provided informed consent in accordance with institutional guidelines. Participants were briefed on the purpose of the study, the modalities collected (RGB video, depth, motion capture, tactile sensing), the intended research use, and their right to withdraw at any time.
Privacy Protection. Although the dataset contains real-world human activities and may capture visible human faces or identifiable biometric patterns, several privacy-preserving strategies were implemented:
Controlled subjects only: No bystanders or non-consenting individuals appear in the dataset; all scenes were recorded in controlled environments with only consenting operators present.
No identity labels: We do not include personal information, demographic information, or identity annotations.
Face and identity protection: When faces are visible in RGB streams, we apply optional face blurring in the public release version.
No audio recordings: We do not collect speech or audio signals to avoid unintended disclosure of personal information.
Sensitive Content and Safety. The dataset does not include minors, hazardous behaviors, or sensitive activities. All recordings were conducted in standard workplace or daily-life environments using safety protocols.
Data Usage and Licensing. The released dataset is intended exclusively for research on embodied AI, manipulation, perception, and VLA learning. Redistribution or commercial use is restricted according to the license accompanying the dataset. Users are required to comply with ethical research practices and avoid any attempts at identity recognition or misuse of human-related data.
Potential Risks and Mitigations. As with any dataset containing human recordings, there is a theoretical risk of model misuse, such as identity extraction or surveillance applications. To mitigate these risks, we provide: (1) no textual identity metadata, (2) optional visual anonymization tools, and (3) a license that explicitly forbids re-identification, face recognition, or surveillance-related uses.
Institutional Review. All data collection procedures were reviewed and approved by an internal ethics committee to ensure compliance with privacy protection and responsible AI principles.